# Brain3D: Generating 3D Objects from fMRI

### v4

@@ Abstract
keywords:
fMRI, 3D generation, diffusion, brain signal decoding
@@ 1 Introduction
Understanding how humans see and perceive the world is a fundamental question in computer vision [3, 4]. The ability to perceive and interpret the environment in three dimensions (3D) is a critical aspect of human vision that enables human to navigate, interact, and make sense of the world around us [5, 6]. Despite significant advancements in the understanding of the neural mechanisms underlying visual perception [7, 8], the intrinsic processes involved in 3D vision remain largely unexplored.
Traditional approaches in neuroscience are limited by ethical concerns [9], interspecies differences [10], and a lack of high-resolution temporal and spatial data, which hinders the translation of findings to human vision. Similarly, while artificial vision models have excelled in computer vision tasks [11, 12, 13, 14], they fall short in mimicking the neural processes of biological vision and fail to incorporate brain signals effectively. Neural decoding models have made strides in reconstructing visual content from fMRI data [15, 16, 17, 18], yet they are constrained by their focus on single views, limiting their ability to fully capture the complexity of human 3D vision.
Consequently, our goal here is to develop a comprehensive computing framework that bridges the gap between artificial vision models and biological vision and gains insights into human visual system, with a specific focus on understanding and simulating the human brain’s 3D visual processing capabilities. To that end, here we explore the intrinsic 3D human vision perception through the lens of fMRI data [19], known for its high spatial resolution, relatively low cost compared to other neuroimaging techniques, and non-invasiveness. We introduce Brain3D, a bio-inspired method capable of understanding how the brain interprets and processes visual stimuli in 3D. Brain3D takes fMRI data of a subject presented with a 2D object image and yields the corresponding 3D object model. There are two key components with Brain3D: tackling the noises with high-level semantic signals by decomposing it into a series of representative embedding vectors, and a two-stage architecture design for progressive high-level information integration. Importantly, our model captures the distinct functionalities and interplay relationships of each region in the human vision system, aligning with established neuroscience discoveries. This alignment allows Brain3D to potentially serve as a useful tool for clinical fMRI evaluation.
@@ 2 Related work
Neuroscience Traditional approaches in neuroscience and life sciences, such as animal studies [20, 21] and invasive techniques [22, 23], have provided valuable insights into neural mechanisms. However, these methods face significant limitations, including the need for animal sacrifice [24], ethical concerns [9], and interspecies heterogeneity [10], which can hinder the translation of findings to human vision. Additionally, traditional methods often lack the spatial and temporal resolution necessary to capture the complex dynamics of neural activity underlying 3D vision perception. Our work offers a novel interdisciplinary approaches that explore intrinsic human vision perception through non-invasive fMRI-based generation.
@@
Computer vision Computer vision, which aims to develop computational models inspired by biological vision [25, 26], has emerged as a promising avenue for understanding the neural basis of visual perception. While vision models have achieved remarkable success in various tasks, such as object recognition [11, 12], detection [13], and scene understanding [14], their contribution to understanding biological vision remains limited. These models primarily focus on catching human-like performance on specific tasks using visual input, without directly incorporating brain signals or mimicking the neural processes involved in biological vision. Furthermore, these models are not designed to simulate the human visual system. In contrast, this work directly integrates brain signals into a visual computing model, enabling the simulation of 3D vision that aims to reflect the neural mechanisms of biological vision.
@@
Neural decoding Neural decoding models have focused on extracting essential representations of brain signals for visual content prediction [27, 28] and object recognition [29]. These investigations have also shown promise in reconstructing images [30] from fMRI data using techniques such as generative adversarial networks [15, 16] and latent-space diffusion [17, 18]. However, these studies are currently limited to single views, which constrains their ability to model and analyze the underlying human brain’s 3D visual processing capabilities. When considering brain-signal-based 3D synthesis, the intrinsically significant diversity within the human brain can cause instability, resulting in a lack of consistency for 3D synthesis. To bridge this gap, our work pioneers the use of fMRI data for generating 3D object models, capturing the complex interplay between different regions of the human visual system.
@@
Cross-modality 3D generation Previous advancements in cross-modality 3D generation are limited in text-based [31, 32, 33, 34] and image-based [35, 36, 37] pipelines. These generation approaches leverage score distillation sampling [31] (SDS) and extensive 3D datasets [38, 39] to distill text or image into 3D representations, including NeRF [40], DMTet [41] or Gaussian splatting [42]. In comparison, our work builds on the concept of hybrid SDS [37] to distill brain signals into both high-level semantics and low-level visuals. This distinctive capability allows our model to achieve cross-modal 3D generation from solely brain signals, even without external images or texts as conditional guidance.
@@ 3.1 Overview
Our key technical contribution lies in the novel design of a UMAP-based projection [43], which effectively addresses high-level semantic noise, allowing us to decompose these embeddings into separate distinct representations. This enhancement is crucial for deriving 3D multi-view semantics from brain signals, thereby facilitating accurate 3D generation. Subsequently, we design a two-stage diffusion method [50, 48, 51] to reconstruct the corresponding 3D objects based soley on brain signals. 3D objects are generated through distilling a 2D diffusion model with condition on both the high level and low level fMRI embeddings, using Score Distillation Sampling (SDS) [47](Fig. 2e). The results affirm Brain3D’s effectiveness in modeling human vision process and diagnosing neurological abnormalities across multiple brain regions, including V1, V2, V3, V4, and MTL.
@@ 3.2 Preliminary: Diffusion based 3D generation
Our approach employs a diffusion model [50] aimed at manipulating data through a two-phased process. Initially, a forward process introduces Gaussian noise with parameter into the data to achieve noised input , which can be mathematically represented as follows:
@@
This mechanism is further extended to 3D generation through the integration of SDS [47]. It involves two key components: a scene model and a pre-trained Text-to-Image (T2I) diffusion model. The scene model, parameterized by , generates differentiable images through . The pre-trained T2I diffusion model predicts the sampled noise conditioned on the noisy image , created by adding Gaussian noise with a random scale factor to . The optimization of this extended model is guided by the following objective:
@@ 3.3 Tackling the noises with high-level semantics
To solve the above problem, we design a novel approach using the Uniform Manifold Approximation and Projection (UMAP) [43] (Fig. 2b). Unlike MindEye using UMAP solely for visualization, we incorporate UMAP projection into our model architecture to effectively manage high-level semantic noise. The key idea is to distill the high-level semantic embeddings into a series of representative embedding vectors with a unique projection.
Formally, following the transformation of fMRI into a high-level embedding by the MLP backbone , this embedding is subsequently projected into a set of stable embeddings, , derived from the Generative Text-to-Image Transformer [52]. In our design, we concatenate the embeddings extracted from all training images. This results in a comprehensive set of embeddings, , for each image. These combined embeddings form a high-dimensional dataset, denoted as D.
@@ 3.4 Two-stage 3D generation
To effectively leverage the projected high-level embeddings, we devise a two-stage 3D generation architecture characterized by progressively integrating high-level information during 3D object modeling(Fig. 2e). The first stage is a perceptual stage, utilizing the Neural Radiance Fields (NeRF) [45] for 3D representation. It defines an implicit 3D representation framework as:
This framework allows the gradients from the Score Distillation Sampling (SDS) [47] to directly influence the output , enabling simultaneous optimization of the density Multilayer Perceptron (MLP) and color MLP . As a result, NeRF is particularly effective at modifying both the color and shape of the 3D object in this stage. In our implementation, we incorporate Zero-1-to-3 [49] for low-level guidance and Stable Diffusion [48] for high-level guidance.
Subsequently, a semantic stage follows where we employ DMTet [41] to convert the above NeRF into a 3D mesh [46]. This conversion prepares the 3D model for further processing. We continue to harness Stable Diffusion [48], coupled with a depth-conditioned Control-Net [53] for high-level guidance. While the 3D mesh does not possess the same versatility as NeRF in shaping modifications, it excels in capturing detailed RGB color information. Therefore, after establishing the coarse shape in the perceptual stage, we shift our focus to refining the color details in this stage. This two-tier design allows for an intricate and nuanced development of the 3D object, leading to a final output that is both elaborately detailed and of high resolution.
@@ 3.5 Training and inference
During training, we optimize the parameters for both low-level encoder and high-level encoder, leaving all the rest components frozen. The high-level encoder is detailed in Sec. 3.3, while the low-level encoder comprises MLP backbone, diffusion prior [54], and Versatile diffusion [44]. We use a training dataset with image-text pairs. Inspired by MindEye [18], the loss function for low-level and high-level embeddings are designed with integration of bidirectional CLIP [1] contrastive learning and MixUp [2] augmentation to improve generalization in the beginning of training stage, as described in the follows:
@@
The in Eq. (13) represents a fMRI embedding derived from either a low-level encoder or a high-level encoder , denoted as , using the initial fMRI . This process is obtained with MixUp augmentation [2] as the subsequent equations:
After one third of the training stage, the BiMixCo loss for both and is replaced with soft contrastive loss to facilitate softmax probability distribution for improved knowledge distillation [55], expressed as:
@@ 4.1 Dataset
We conduct extensive experiments on two publicly available fMRI datasets: the Natural Scenes Dataset (NSD) [19] and the Generic Object Decoding (GOD) [56] dataset. NSD [19] is comprised of high-resolution whole-brain fMRI measurements (1.8-mm isotropic resolution, 1.6-s sampling rate) from eight healthy adult subjects. These subjects were observed while viewing thousands of 2D colored natural images sampled from COCO dataset [57], spread across 30 to 40 scanning sessions. The two-stage training set includes 8,859 images and 24,980 fMRI trials, whereas the test set contains 982 images and 2,770 fMRI trials. To enhance data quality, we compute average fMRI responses for images that are presented more than once. GOD [56] consists of fMRI recordings from five subjects viewing ImageNet stimuli, resulting in a training set of 1,200 images (8 images per 150 categories) and a test set of 50 images (1 from each of 50 categories). Each subject’s data amounts to 1,200 training and 1,750 testing fMRI trials. While offering no comprehensive annotations for ROI-based analysis as NSD, this dataset enables generalization and robustness evaluation of our method across different stimuli and scanning conditions.
@@ 4.2 Evaluation metrics
To analyze the generation quality, we introduce four fMRI-to-3D evaluation metrics: 3D Inception [61], 3D CLIP [62], 3D EfficientNet [63], and 3D Swav [64]. These metrics are designed to overcome the limitations of traditional pixel-by-pixel image reconstruction metrics like PixelCorrelation [65], which lacks robustness against rotations and translations in 3D space, and are built upon well-established metrics in previous neural decoding researches [17, 18] to ensure reliability. For comprehensive evaluation, we construct a level-2 icosahedron with a 2.2 radius around the generated 3D object and capture the 3D object from all the icosahedron vertices to achieve multi-view coverage, similar to the approach in the text-to-3D benchmark [66]. The metrics 3D Inception and 3D CLIP, utilizing the final pooling layer of InceptionV3 [61] and CLIP model [62] respectively, assess the accuracy of two-way identification between original image embeddings and corresponding brain embeddings. Consequently, larger 3D Inception and 3D CLIP represent higher matching accuracy in features between ground truth images and generated 3D objects, indicating a well extracted features from fMRI.
On the other hand, 3D EfficientNet and 3D Swav employ EfficientNet-B1 [63] and SwAV-ResNet50 [64] to calculate the average correlation distance between the extracted features. Lower correlation distance represents higher correlation between ground truth images and generated 3D objects, representing better reconstruction.
@@ 4.3 Qualitative evaluation
To assess the generalizability and practical utility in clinical settings of Brain3D, we perform comprehensive comparisons with previous 2D methods [18, 58, 59, 60]. We could not perform quantitative comparison as all previous 2D methods do not output 3D generation as ours. As shown in Fig. 4, from the fMRI data Brain3D extracts rich appearance, semantic, and geometric information, presenting a high degree of consistency with the stimuli images. In the left columns of Fig. 4a, the generated swan and lizard from Brain3D not only show the highest semantic and texture consistency with the stimulus, but also preserve 3D geometry details inferred from brain signals. Despite limited observation with the giraffe, skier and teddy bear in the three cases (see the last two rows of Fig. 4b), our bio-inspired model can process 3D details and comprehend their semantics beyond what is immediately visible in the 2D image, achieving finer semantic details than MindEye [18]. As shown in Fig. 5c with both hemispheres, our model also achieves a significant quantitative neural decoding performance across all samples in the testing set. This result underscores the remarkable human’s neural processing capability in constructing intricate 3D structures by interpreting and internalizing 3D geometric information from 2D visual stimuli [5, 6].

### v5

@@ 1 Introduction
Understanding how humans see and perceive the world is a fundamental question in computer vision Hubel and Wiesel (1979); Hupé et al. (1998). The ability to perceive and interpret the environment in three dimensions (3D) is a critical aspect of human vision that enables human to navigate, interact, and make sense of the world around us Georgieva et al. (2009); Groen and Baker (2019). Despite significant advancements in the understanding of the neural mechanisms underlying visual perception Tong (2003); Popham et al. (2021), the intrinsic processes involved in 3D vision remain largely unexplored.
Traditional approaches in neuroscience are limited by ethical concerns Díaz et al. (2021), interspecies differences Jansen et al. (2020), and a lack of high-resolution temporal and spatial data, which hinders the translation of findings to human vision. Similarly, while artificial vision models have excelled in computer vision tasks Kubilius et al. (2019); Sexton and Love (2022); Baek et al. (2021); Frey et al. (2023), they fall short in mimicking the neural processes of biological vision and fail to incorporate brain signals effectively. Neural decoding models have made strides in reconstructing visual content from fMRI data Seeliger et al. (2018); Shen et al. (2019); Chen et al. (2023a); Scotti et al. (2024), yet they are constrained by their focus on single views, limiting their ability to fully capture the complexity of human 3D vision.
Consequently, our goal here is to develop a comprehensive computing framework that bridges the gap between artificial vision models and biological vision and gains insights into human visual system, with a specific focus on understanding and simulating the human brain’s 3D visual processing capabilities. To that end, here we explore the intrinsic 3D human vision perception through the lens of fMRI data Allen et al. (2022), known for its high spatial resolution, relatively low cost compared to other neuroimaging techniques, and non-invasiveness. We introduce Brain3D, a bio-inspired method capable of understanding how the brain interprets and processes visual stimuli in 3D. Brain3D takes fMRI data of a subject presented with a 2D object image and yields the corresponding 3D object model. There are two key components with Brain3D: tackling the noises with high-level semantic signals by decomposing it into a series of representative embedding vectors, and a two-stage architecture design for progressive high-level information integration. Importantly, our model captures the distinct functionalities and interplay relationships of each region in the human vision system, aligning with established neuroscience discoveries. This alignment allows Brain3D to potentially serve as a useful tool for clinical fMRI evaluation.
@@ 2 Related work
Neuroscience Traditional approaches in neuroscience and life sciences, such as animal studies Costa et al. (2022); Chen and Rechavi (2022) and invasive techniques Singh et al. (2022); Risom et al. (2022), have provided valuable insights into neural mechanisms. However, these methods face significant limitations, including the need for animal sacrifice Rai and Kaushik (2018), ethical concerns Díaz et al. (2021), and interspecies heterogeneity Jansen et al. (2020), which can hinder the translation of findings to human vision. Additionally, traditional methods often lack the spatial and temporal resolution necessary to capture the complex dynamics of neural activity underlying 3D vision perception. Our work offers a novel interdisciplinary approaches that explore intrinsic human vision perception through non-invasive fMRI-based generation.
@@
Computer vision Computer vision, which aims to develop computational models inspired by biological vision Khaligh-Razavi and Kriegeskorte (2014); LeCun et al. (2015), has emerged as a promising avenue for understanding the neural basis of visual perception. While vision models have achieved remarkable success in various tasks, such as object recognition Kubilius et al. (2019); Sexton and Love (2022), detection Baek et al. (2021), and scene understanding Frey et al. (2023), their contribution to understanding biological vision remains limited. These models primarily focus on catching human-like performance on specific tasks using visual input, without directly incorporating brain signals or mimicking the neural processes involved in biological vision. Furthermore, these models are not designed to simulate the human visual system. In contrast, this work directly integrates brain signals into a visual computing model, enabling the simulation of 3D vision that aims to reflect the neural mechanisms of biological vision.
@@
Neural decoding Neural decoding models have focused on extracting essential representations of brain signals for visual content prediction Naselaris et al. (2011); Kamitani and Tong (2005) and object recognition Horikawa and Kamitani (2017a). These investigations have also shown promise in reconstructing images Beliy et al. (2019) from fMRI data using techniques such as generative adversarial networks Seeliger et al. (2018); Shen et al. (2019) and latent-space diffusion Chen et al. (2023a); Scotti et al. (2024). However, these studies are currently limited to single views, which constrains their ability to model and analyze the underlying human brain’s 3D visual processing capabilities. When considering brain-signal-based 3D synthesis, the intrinsically significant diversity within the human brain can cause instability, resulting in a lack of consistency for 3D synthesis. To bridge this gap, our work pioneers the use of fMRI data for generating 3D object models, capturing the complex interplay between different regions of the human visual system.
@@
Cross-modality 3D generation Previous advancements in cross-modality 3D generation are limited in text-based (Poole et al., 2022; Lin et al., 2023; Wang et al., 2023; Tang et al., 2024) and image-based (Liu et al., 2023a; Shi et al., 2023; Qian et al., 2023) pipelines. These generation approaches leverage score distillation sampling (Poole et al., 2022) (SDS) and extensive 3D datasets (Deitke et al., 2023; Yu et al., 2023) to distill text or image into 3D representations, including NeRF (Mildenhall et al., 2020), DMTet (Shen et al., 2021) or Gaussian splatting (Kerbl et al., 2023). In comparison, our work builds on the concept of hybrid SDS Qian et al. (2023) to distill brain signals into both high-level semantics and low-level visuals. This distinctive capability allows our model to achieve cross-modal 3D generation from solely brain signals, even without external images or texts as conditional guidance.
@@ 3.1 Overview
Our key technical contribution lies in the novel design of a UMAP-based projection McInnes et al. (2018), which effectively addresses high-level semantic noise, allowing us to decompose these embeddings into separate distinct representations. This enhancement is crucial for deriving 3D multi-view semantics from brain signals, thereby facilitating accurate 3D generation. Subsequently, we design a two-stage diffusion method Ho et al. (2020); Rombach et al. (2022); Saharia et al. (2022) to reconstruct the corresponding 3D objects based soley on brain signals. 3D objects are generated through distilling a 2D diffusion model with condition on both the high level and low level fMRI embeddings, using Score Distillation Sampling (SDS) Poole et al. (2022)(Fig. 2e). The results affirm Brain3D’s effectiveness in modeling human vision process and diagnosing neurological abnormalities across multiple brain regions, including V1, V2, V3, V4, and MTL.
@@ 3.2 Preliminary: Diffusion based 3D generation
Our approach employs a diffusion model Ho et al. (2020) aimed at manipulating data through a two-phased process. Initially, a forward process introduces Gaussian noise with parameter into the data to achieve noised input , which can be mathematically represented as follows:
@@
This mechanism is further extended to 3D generation through the integration of SDS Poole et al. (2022). It involves two key components: a scene model and a pre-trained Text-to-Image (T2I) diffusion model. The scene model, parameterized by , generates differentiable images through . The pre-trained T2I diffusion model predicts the sampled noise conditioned on the noisy image , created by adding Gaussian noise with a random scale factor to . The optimization of this extended model is guided by the following objective:
@@ 3.3 Tackling the noises with high-level semantics
To solve the above problem, we design a novel approach using the Uniform Manifold Approximation and Projection (UMAP) McInnes et al. (2018) (Fig. 2b). Unlike MindEye using UMAP solely for visualization, we incorporate UMAP projection into our model architecture to effectively manage high-level semantic noise. The key idea is to distill the high-level semantic embeddings into a series of representative embedding vectors with a unique projection.
Formally, following the transformation of fMRI into a high-level embedding by the MLP backbone , this embedding is subsequently projected into a set of stable embeddings, , derived from the Generative Text-to-Image Transformer Wang et al. (2022). In our design, we concatenate the embeddings extracted from all training images. This results in a comprehensive set of embeddings, , for each image. These combined embeddings form a high-dimensional dataset, denoted as D.
@@ 3.4 Two-stage 3D generation
To effectively leverage the projected high-level embeddings, we devise a two-stage 3D generation architecture characterized by progressively integrating high-level information during 3D object modeling(Fig. 2e). The first stage is a perceptual stage, utilizing the Neural Radiance Fields (NeRF) Mildenhall et al. (2020) for 3D representation. It defines an implicit 3D representation framework as:
This framework allows the gradients from the Score Distillation Sampling (SDS) Poole et al. (2022) to directly influence the output , enabling simultaneous optimization of the density Multilayer Perceptron (MLP) and color MLP . As a result, NeRF is particularly effective at modifying both the color and shape of the 3D object in this stage. In our implementation, we incorporate Zero-1-to-3 Liu et al. (2023b) for low-level guidance and Stable Diffusion Rombach et al. (2022) for high-level guidance.
Subsequently, a semantic stage follows where we employ DMTet Shen et al. (2021) to convert the above NeRF into a 3D mesh Hoppe et al. (1993). This conversion prepares the 3D model for further processing. We continue to harness Stable Diffusion Rombach et al. (2022), coupled with a depth-conditioned Control-Net Zhang et al. (2023) for high-level guidance. While the 3D mesh does not possess the same versatility as NeRF in shaping modifications, it excels in capturing detailed RGB color information. Therefore, after establishing the coarse shape in the perceptual stage, we shift our focus to refining the color details in this stage. This two-tier design allows for an intricate and nuanced development of the 3D object, leading to a final output that is both elaborately detailed and of high resolution.
@@ 3.5 Training and inference
During training, we optimize the parameters for both low-level encoder and high-level encoder, leaving all the rest components frozen. The high-level encoder is detailed in Sec. 3.3, while the low-level encoder comprises MLP backbone, diffusion prior Ramesh et al. (2022), and Versatile diffusion Xu et al. (2023). We use a training dataset with image-text pairs. Inspired by MindEye Scotti et al. (2024), the loss function for low-level and high-level embeddings are designed with integration of bidirectional CLIP Radford et al. (2021a) contrastive learning and MixUp Zhang et al. (2017) augmentation to improve generalization in the beginning of training stage, as described in the follows:
@@
The in Eq. (13) represents a fMRI embedding derived from either a low-level encoder or a high-level encoder , denoted as , using the initial fMRI . This process is obtained with MixUp augmentation Zhang et al. (2017) as the subsequent equations:
After one third of the training stage, the BiMixCo loss for both and is replaced with soft contrastive loss to facilitate softmax probability distribution for improved knowledge distillation Hinton et al. (2015), expressed as:
@@ 4.1 Dataset
We conduct extensive experiments on two publicly available fMRI datasets: the Natural Scenes Dataset (NSD) Allen et al. (2022) and the Generic Object Decoding (GOD) Horikawa and Kamitani (2017b) dataset. NSD Allen et al. (2022) is comprised of high-resolution whole-brain fMRI measurements (1.8-mm isotropic resolution, 1.6-s sampling rate) from eight healthy adult subjects. These subjects were observed while viewing thousands of 2D colored natural images sampled from COCO dataset Lin et al. (2014), spread across 30 to 40 scanning sessions. The two-stage training set includes 8,859 images and 24,980 fMRI trials, whereas the test set contains 982 images and 2,770 fMRI trials. To enhance data quality, we compute average fMRI responses for images that are presented more than once. GOD Horikawa and Kamitani (2017b) consists of fMRI recordings from five subjects viewing ImageNet stimuli, resulting in a training set of 1,200 images (8 images per 150 categories) and a test set of 50 images (1 from each of 50 categories). Each subject’s data amounts to 1,200 training and 1,750 testing fMRI trials. While offering no comprehensive annotations for ROI-based analysis as NSD, this dataset enables generalization and robustness evaluation of our method across different stimuli and scanning conditions.
@@ 4.2 Evaluation metrics
To analyze the generation quality, we introduce four fMRI-to-3D evaluation metrics: 3D Inception Szegedy et al. (2016), 3D CLIP Radford et al. (2021b), 3D EfficientNet Tan and Le (2019), and 3D Swav Caron et al. (2020). These metrics are designed to overcome the limitations of traditional pixel-by-pixel image reconstruction metrics like PixelCorrelation Potetz and Lee (2003), which lacks robustness against rotations and translations in 3D space, and are built upon well-established metrics in previous neural decoding researches Chen et al. (2023a); Scotti et al. (2024) to ensure reliability. For comprehensive evaluation, we construct a level-2 icosahedron with a 2.2 radius around the generated 3D object and capture the 3D object from all the icosahedron vertices to achieve multi-view coverage, similar to the approach in the text-to-3D benchmark He et al. (2023). The metrics 3D Inception and 3D CLIP, utilizing the final pooling layer of InceptionV3 Szegedy et al. (2016) and CLIP model Radford et al. (2021b) respectively, assess the accuracy of two-way identification between original image embeddings and corresponding brain embeddings. Consequently, larger 3D Inception and 3D CLIP represent higher matching accuracy in features between ground truth images and generated 3D objects, indicating a well extracted features from fMRI.
On the other hand, 3D EfficientNet and 3D Swav employ EfficientNet-B1 Tan and Le (2019) and SwAV-ResNet50 Caron et al. (2020) to calculate the average correlation distance between the extracted features. Lower correlation distance represents higher correlation between ground truth images and generated 3D objects, representing better reconstruction.
@@ 4.3 Qualitative evaluation
To assess the generalizability and practical utility in clinical settings of Brain3D, we perform comprehensive comparisons with previous 2D methods Scotti et al. (2024); Chen et al. (2023b); Ozcelik et al. (2022); Gaziv et al. (2022). We could not perform quantitative comparison as all previous 2D methods do not output 3D generation as ours. As shown in Fig. 4, from the fMRI data Brain3D extracts rich appearance, semantic, and geometric information, presenting a high degree of consistency with the stimuli images. In the left columns of Fig. 4a, the generated swan and lizard from Brain3D not only show the highest semantic and texture consistency with the stimulus, but also preserve 3D geometry details inferred from brain signals. Despite limited observation with the giraffe, skier and teddy bear in the three cases (see the last two rows of Fig. 4b), our bio-inspired model can process 3D details and comprehend their semantics beyond what is immediately visible in the 2D image, achieving finer semantic details than MindEye Scotti et al. (2024). As shown in Fig. 5c with both hemispheres, our model also achieves a significant quantitative neural decoding performance across all samples in the testing set. This result underscores the remarkable human’s neural processing capability in constructing intricate 3D structures by interpreting and internalizing 3D geometric information from 2D visual stimuli Georgieva et al. (2009); Groen and Baker (2019).